write the following equations in standard form:
1.4x^2+4y^2+8x-24y=60
2.4x^2-6y^2+8x-24y=68

Answer by lwsshak3(11628)

write the following equations in standard form:
1.4x^2+4y^2+8x-24y=60
Complete the square
4(x^2+2x+1)+4(y^2-6y+9)=60+36+4
4(x+1)^2+4(y-3)^2=100
divide by 4
(x+1)^2+(y-3)^2=25
This is an equation of a circle with center at (-1, 3) and radius=5
..
2. 4x^2-6y^2+8x-24y=68
Complete the square
4(x^2+2x+1)-6(y^2+4y+4)=68+4-24
4(x+1)^2-6(y+2)^2=48
divide by 48
(x+1)^2/12-(y+2)^2/8=1
This is an equation of a hyperbola with horizontal transverse axis and center at (-1,-2)